Switched to W1 electronics box, because W2 was continuously resetting the BCU39 board.

Mirror configuration

NOTE: ICC is active (control current integrator)

Remind that the mirror has 19 not-working actuators: #modes = #working-actuators = 653

For this particular mesurement the actuators are numebered in a special way with respect the standard mirror or dsp ordering. Only the working actuators are considered and are ordered following the "mirror" ordering (stepping over the not working actuators) from 0 (inner ring) to 652 (outer ring)

HEXAPOD (with tower cooling on: [-552.108, 2231.62, -116.872, 165.842, -45.7826, -0.060851]=[x,y,z,a,b,c])

WFS configuration:

  • Select "Power on config": Left_Box
  • Press "Power On" button and wait for the end of the power-on process
  • Set:
    • Binning to 1
    • Loop frequency to 25Hz (IMPORTANT for interferometer setting)
      • With 25Hz the CCD39 is set with a frame readout of 150kpix/sec (= readout in 42.7ms)
    • Modulation to 0.0
    • Press button "Apply" to apply above settings and wait for AOSet status

4D configuration:

  • From "Camera Settings":
    • External Trigger: TTL0
    • Delay: 8ms
  • New setting (no used before): Data Fill=10 (interpolation of missing data)
  • exit 4Sight and start again for enable the trigger settings
  • the frames should be updated, otherwise some error in the connections between W and 4D trigger would be present
  • In case some disturb appear on the top-left of the frame due to the frame grabber (static pattern): open the "Camera Settings" panel and slide back and forth the "Delay" slider. The pattern should disappear.(ATTENTION: this operation is no longer allowed: later we discovered that this operation screws up the triggering frequency)
  • Run the default script by pressing the green "D>" button
  • Select from menu: Tools->Debug->Open 4Sight Scripting Console
  • Execute "from I4D import comm4d" command in the Scripting Console

Tower configuration:

  • Switch off the Tower cooling by setting 0 to the "Working Mode" to 0
  • The hexapodhas to be moved to remove tilt introducedby switching off the tower cooling pump
  • The Temperature settings from the evening befor (running all the night):
    • Both Bench and Mirror: 20.0C (SRC) 19.8C (Valve)

Sequence of Operations:

This first operations will not produce the final measurements, the aim is to have a first set of data to give to Fernando. For instance the inital flattening of the mirror is not optimal and modal amplitudes could not be tuned.

Test of noise level running measurements without disturbance.

  • On W
    • Be sure "Disturbance" is "Disabled"
    • Close the loop
  • On 4D scripting Console
    • comm4d.capture(300, "2009_08_08_noise", "2009_08_08")
      • to capture 300 measures having filename 2009_08_08_noise_nnnn.dat in the sub-directory 2009_08_08\raw of the ram disk (Z:\4D\)
  • On W
    • Open the loop
  • On 4D scripting Console
    • comm4d.move("2009_08_08")
      • to move the 2009_08_08 folder from the ram disk to the D:\Zcopy directory in the hard-disk
    • comm4d.process(300,"2009_08_08")
      • to pre-process (calibration) of the 300 measurament from 2009_08_08\raw directory and save the results in 2009_08_08\calibrated (same name and extension)
    • comm4d.reduce(300, "2009_08_08_noise", "2009_08_08")
      • to save the calibrated 300 measurements from 2009_08_08\calibrated in DHF5 format in the directory 2009_08_08\dhf5 (same name, but different extension: .dh5)
  • Measurement of first 100 modes IF for Fernando's software check

  • On AdSec
    • print, make_modal_disturb(0,99,/SEND)
      • to build the disturbance history of 4000 (fixed number) samples, the first 300 steps corespond to up-down-up (3 steps per mode) sequenceof the 100 modes (from 0 to 99, see commando line). The /SEND keyword transfers the sequence on the secondary mirror DSP memory to be used. The disturbance history is padded with zeros to obtained the 4000 steps sequence.
  • On W
    • Be sure "Disturbance" is "Disabled"
    • Close the loop
  • On 4D
    • type on Scripting Console (WITHOUT EXECUTING) the command line:
    • comm4d.capture(500, "m000-099", "2009_08_08_m000-099_a")
      • to grab 500 frames (300 are needed for 100 modes + 200/25Hz=8sec for manual sinc buffer and noise measurement
  • On W
    • Point the mouse on the "Sync WFS" button
  • Hit return-key on 4D Scripting Consol to start the acquisition AND about 1-2sec later press the W "Sync WFS" to start the disturbance
  • when the acquisition on 4D ends and no disturbance deformation should be seen on the interferometer, DISABLE the disturbance on W.
    • It is important to disablethe disturbance, because it is applied cyclicaly
  • in case the user will post-process the data later, the loop can be kept closed (with disturbance disabled, a new disturbance can be configured with the make_modal_disturb IDL command and restart from the comm4d.capture command on the 4D scripting console
  • in case the post-process is required immediatly after the capture, process the data as in the previous section (using the comm4d commands)
ATTENTION: after post-processing this set of data shows no modes applied. Use these data for noise evaluation

Measurement of zonal IF

Same procedure as above. The sequence for the zonal IF is sent to the mirror with the following command:

print, make_modal_disturb(first_act_num,last_act_num,/ZONAL,/SEND)

Because we have a limitation of 1700 frames on the 4D (Ram Disk memory) and we need 672*3=2013 frames, we will take the data in to sets of 336 actuators:
  • First set of zonal IF (from act num 0 to 335):
    • directory: 2009_08_08_z000-335_a
    • filename: z000-335_nnnn.h5
  • Second set of IF (from act num 336 to 671):
    • directory: 2009_08_08_z336-671_a
    • filename: z336-671_nnnn.h5

Measurement of noise levels

  • With TOWER cooling ON:
    • directory: 2009_08_08_noise_tower_on_a
    • filename: noise_tower_on_nnnn.h5
    • number of 4D mesurements 500
  • With TOWER cooling OFF:
    • directory: 2009_08_08_noise_tower_off_a
    • filename: noise_tower_off_nnnn.h5
    • number of 4D mesurements 500
  • With MIRROR and TOWER cooling OFF:
    • directory: 2009_08_08_noise_mirror_off_a
    • filename: noise_mirror_off_nnnn.h5
    • number of 4D mesurements 500

The cooling of the mirror is switched on again after the last measurement and the tower cooling has been kept off

ERROR!

Looking at the log of the 4D, the acquisition frequency was always wrong (about 12.5Hz instead of 25Hz, even if the WFS was set and running properly)

ALL THE MEASUREMENT HAVE TO BE REDONE

Measurement of noise levels (set: "_b")

  • With TOWER cooling ON:
    • directory: 2009_08_08_noise_tower_on_b
    • filename: noise_tower_on_nnnn.h5
    • number of 4D mesurements 500
  • With TOWER cooling OFF:
    • directory: 2009_08_08_noise_tower_off_b
    • filename: noise_tower_off_nnnn.h5
    • number of 4D mesurements 500
  • With MIRROR and TOWER cooling OFF:
    • directory: 2009_08_08_noise_mirror_off_b
    • filename: noise_mirror_off_nnnn.h5
    • number of 4D mesurements 500

The cooling of the mirror and the tower cooling has been kept off to perform the following measurement

Measurement of zonal IF with AdSec cooling switched off (Set: "_b")

  • First set of zonal IF (from act num 0 to 335):
    • directory: 2009_08_08_z000-335_b
    • filename: z000-335_nnnn.h5
    • number of 4D mesurements 1200
  • Second set of IF (from act num 336 to 671):
    • directory: 2009_08_08_z336-671_b
    • filename: z336-671_nnnn.h5
    • number of 4D mesurements 1200

The cooling of the mirror has been restarted after the last measurement.

NOTE: the temperature of the cooling inside the AdSec while the pump was off rised from 24C to 32C. Temperature drift could affect this measurement.

Problems after

When the pump has been restarted the high temperature coolant inside the AdSec has been sensed by the outlet temperature probe (32C) giving a HouseKeeper failure of the mirror (auto-RIP)

Waiting for temeperature stabilization before setting again the shell.

Looking at the mirror shape in xadsec a large gap dip is seen around act 599 (95um). The dip seems to be physical because of the shape correlated with neighbor actuators.
  • Status saved in ./meas/2009_08_08_possible_contamination.sav

It seems to reducing with time.

We try to apply a power-off/power-on cycle to the mirror.

The power cycling made the effect disappear. It was not a contamination, just an effect of the electronics due to the fail stop of the overheating.

Temperature behavior of the various parts are back to normal values.

Restart the measurement process.

Start raining

Measurement of zonal IF (Set: "_c")

  • First set of zonal IF (from act num 0 to 335):
    • directory: 2009_08_08_z000-335_c
    • filename: z000-335_nnnn.h5
    • number of 4D mesurements 1200
      • first actuation of act 8 is in frame 46 (identified by jump from ring 0 of act 8 to ring 1 of act 9)
      • only 8 acts in ring 0 (one is missing, because not working), so:
        • act 8 frame 46
        • act 7 frame 43
        • act 6 is missing because not working
        • act 5 frame 40
        • act 4 frame 37
        • act 3 frame 34
        • act 2 frame 31
        • act 1 frame 28
        • act 0 frame has the first frame without actuation: 25 has no actuation (should have +1), 26 has -1, 27 has +1
        • POSTPROCESSING SHOWED THAT THE FIRST STEP OF THE DISTURBANCE IS NOT APPLIED. ADD ZEROS AT THE BEGINNING OF THE SEQUENCE TO PATCH THIS BUG
  • Second set of IF (from act num 336 to 671):
    • directory: 2009_08_08_z336-671_c
    • filename: z336-671_nnnn.h5
    • number of 4D mesurements 1200
      • in this set the first actuation is in frame 0028
      • first actuation of act 337 is 0030
      • first actuation of the disturbance (first actuation of act 336) has not been applied. Act 336 has not +1,-1,+1 actuations, but only -1(frame 0028), +1(frame0029). Same bug as before: it is a sistematic problem of the disturbance application in the AsSec

Measurement of modal IF (Set: "_d")

  • First set of zonal IF (from act num 0 to 335):
    • directory: 2009_08_08_m000-335_d
    • filename: m000-335_nnnn.h5
    • number of 4D mesurements 1200
      • in this set the first actuaton is in frame 0025
  • Second set of IF (from act num 336 to 671):
    • directory: 2009_08_08_m336-671_d
    • filename: m336-671_nnnn.h5
    • number of 4D mesurements 1200

18:45 end of session. Thunders.

NOTES AND REQUESTS OF SOFTWARE CHANGE:

  • the IDL reduction program (if_redux) is not able to correctly identify the first actuated frame when zonal influence functions are used (the opd rms has small variations with zonal actuations). Particular problematic when zonal actuation is starting from act 0 (in the center with partial masking of pupil). POSSIBLE SOLUTION: a known pattern could be used as first actuation to simplify the identification of the first frame (large astig?)
  • the startup of the disturbance seems to have some problems at the start loosing the first step of the disturbance. Add few zeros (at least one, 5 is better) at the beginning of the command sequence.(5zeri+trifoil+3zeri)
  • add the printing on the "4D scripting console" of the measured acquisition frequency, in order to not open every time the log to read it
  • the commanded amplitude has to be saved together the interferometric data after post processing to relate commands and optical signals. For the modal history toghether with the actuator based commands, also the modal amplitude sequence has to be passed and saved (nel mode_XXX mettici il comando+adsec_save)
  • in the zonal history, if a larger number of working actuators is requested (es 672 when some of the acts are out-of-order) the sequence of the last working act is repeated to pad the missing actuations (forza a zero quelli che non ci sono)
  • The sequence of [+1,-1,etc] has to be saved with the disturbance history in order to simplify the post porcessing of the interferometer data
  • Try taking measurement without ICC
  • Flattening has to be refined
  • l'ordinamento delle IF zonali non e' ne' DSP, ne' mirror (simile al mirror, ma saltando gli attuatori non funzionanti). Questo incsina il post processing introducendo una terza numerazione. Usare mirror puro con attuazione zero degli attuatori non funzionanti o righe programma che facciano passare da una numerazione all'altra.

Attachments

2009_08_08_4D-log.out: 4D Log file

-- ArmandoRiccardi - 2009-08-08
Topic revision: r9 - 21 Aug 2009, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback